OpenStudy (anonymous):

root3 SINA - COSA=0 AND 0 DEGREE IS LESS THAN A AND is less than 90 degree, find the value of A

OpenStudy (anonymous):

\[\sqrt{3}\sin \theta - \cos \theta =0 and 0<\theta <90\]